# A Survey on LLM-as-a-Judge

### v1

@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024c) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Chen, 2023) further codifies this form-filling paradigm, efficiently output evaluation results in JSON dictionary format and obtain multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
@@ 3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023d) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generate responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024b) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
@@ 3.3.1. Integration of Multiple Evaluation Results
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and the integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023d) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022a) and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2024b) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
@@ 3.3.2. Direct Optimization of LLMs’ Outputs
However, methods like score smoothing, which combine implicit logits and explicit outputs require the LLMs to be open-source, or to provide interfaces that allow access to token probabilities, which brings some limitations. Inspired by the work of Weng et al.(Weng et al., 2023) and Madaan et al.(Madaan et al., 2024), self-verification can be used to filter out the evaluation results without sufficient robustness. For example, TrueTeacher(Gekhman et al., 2023) applies self-verification in its evaluation of distilled data by asking the LLM evaluator for its certainty about the evaluation results after providing them, and retaining only those results that pass self-verification. Self-verification is suitable for all LLMs and require no complex computing and processing.
@@ 4. Evaluation of LLM Evaluators
Despite their impressive performance, LLMs exhibit several notable shortcomings, such as hallucinations (Tonmoy et al., 2024), biases (Gallegos et al., 2024), and a lack of robustness (Zhu et al., 2023d). When LLMs are employed as evaluators, these inherent issues can lead to suboptimal evaluation outcomes. Therefore, it is crucial to accurately and comprehensively assess the quality of LLM-as-a-judge and identify potential vulnerabilities. This section will review existing work on the evaluation of LLM-as-a-judge, focusing on three key areas: base metric (Section 4.1), bias (Section 4.2), and robustness (Section 4.3).
@@ 4.1. Basic Metric
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ is the evaluation result of LLM evaluator and human judge respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024c; Bai et al., 2024b) are also employed to access agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision, recall, and F1 scores to evaluate the performance (Zhu et al., 2023b; Wang et al., 2023d).
Both of above metrics rely on the datasets with LLM-generated response and responding human judgements. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. In (Zheng et al., 2023d), MTBench and Chatbot Arena Conversations are proposed. The former has only 80 human-crafted queries, each with several LLMs’ responses and expert-level human annotation on pairwise comparison; the latter is a large collection of crowdsourced data, with more than 30K queries from real-world users and their vote on pairs of responses from different LLMs. FairEval (Wang et al., 2023b) is based on the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated labels between ChatGPT and Vicuna responses. PandaLM (Wang et al., 2023d) constructs a test set comprising 999 pairwise samples, with queries from 252 user-oriented instructions in (Wang et al., 2022). LLMEval2 (Zhang et al., 2023d) is much larger than the previous two, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Shepherd (Wang et al., 2023c) collects 352 samples from multiple sources for its critique model as a test set to evaluate the quality of the critiques. Table 1 shows the benchmarks and their statistics.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023d) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023d) 2023 30k Pairwise 3 FairEval (Wang et al., 2023b) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12
@@ 4.2. Bias
In this section, we systematically review various types of biases in the LLM-as-a-judge context, including their definitions, relevant metrics, and datasets that can be used for evaluation.
@@
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024a; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024a). To measure this bias, (Shi et al., 2024) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor response in certain positions. (Wang et al., 2024a) introduced a metric Conflict Rate to measure the percent of disagreement after change the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
@@
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023d; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original response into a more verbose one (Ye et al., 2024; Zheng et al., 2023d). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
@@
Self-Enhancement Bias describe the phenomenon that LLM evaluators may prefer response generated by themselves (Zheng et al., 2023d; Ye et al., 2024). Considering the significant self-enhancement bias, (Ye et al., 2024) suggested that we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
@@
Other Bias. Diversity Bias refers to bias against certain demographic groups (Ye et al., 2024), including certain genders (Chen et al., 2024a), race, and sexual orientation (Kumar et al., 2024b). (Chen et al., 2024a) revealed evaluators’ tendency toward visually appealing content, regardless of its actual validity, such as the text with emoji. Concreteness bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024). Furthermore, LLM evaluators may favor response with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as seyiment bias (Ye et al., 2024; Li and Sinnamon, 2023).
@@
To advance the development of LLM-as-a-Judge systems, future efforts should address two key challenges: (i) Need for Systematic Benchmark. Due to the diversity of biases, it is crucial to propose a systematic benchmark to evaluate the extent of various biases. As shown in Table 1, (Park et al., 2024) proposed EVALBIASBENCH as a test set to measure six types of bias. (Ye et al., 2024) is dedicated to proposing a unified bias testing process, including automated perturbation and a unified metric. They constructed a bias quantification framework CALM, which covers 12 types of bias. Despite these efforts, there is still no systematic benchmark and dataset that includes all types of biases. (ii) Challenges of Controlled Study. When conducting an investigation into a certain type of bias, it is challenging to isolate the specific direction of interest from other biases and quality-related characteristics. For instance, in the case of position bias, lengthening the response could potentially alter the style, fluency, and coherence, or even introduce new biases such as self-enhancement bias. Additionally, the tendency for GPT-4 to favor its own responses over those of GPT-3.5 can be interpreted as either self-enhancement bias or a proper tendency towards higher quality text. Therefore, it is essential for analytical work to carefully control for these variances.
@@ 4.3. Adversarial Robustness
(Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Furthermore, (Zheng et al., 2024) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
5. Meta-evaluation Benchmark
In Section 3, We have introduced improvement strategies adopted by researchers in existing LLM-as-a-judge works to improve the evaluation capabilities of LLMs. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, as shown in Table 1, there is still a lack of meta-evaluation on whether these improvement strategies effectively optimize the LLM evaluators and which dimensions of evaluation performance are being enhanced. It is possible that some improvement strategies fail to enhance the LLM evaluators’ performance or mitigate biases in practical use, leading to computing waste. In this section, based on benchmarks mentioned in Section 4, we organized a robust and scalable meta-evaluation pipeline and benchmark, and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.

### v2

@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024d) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further codifies this form-filling paradigm, efficiently output evaluation results in JSON dictionary format and obtain multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
@@ 3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023d) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generate responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024c) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
@@ 3.3.1. Integration of Multiple Evaluation Results
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and the integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023e) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022) and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2023b) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
@@ 3.3.2. Direct Optimization of LLMs’ Outputs
However, methods like score smoothing, which combine implicit logits and explicit outputs require the LLMs to be open-source, or to provide interfaces that allow access to token probabilities, which brings some limitations. Inspired by the work of Weng et al.(Weng et al., 2023) and Madaan et al.(Madaan et al., 2023), self-verification can be used to filter out the evaluation results without sufficient robustness. For example, TrueTeacher(Gekhman et al., 2023) applies self-verification in its evaluation of distilled data by asking the LLM evaluator for its certainty about the evaluation results after providing them, and retaining only those results that pass self-verification. Self-verification is suitable for all LLMs and require no complex computing and processing.
@@ 4. Evaluation of LLM Evaluators
Despite their impressive performance, LLMs exhibit several notable shortcomings, such as hallucinations (Tonmoy et al., 2024), biases (Gallegos et al., 2024a), and a lack of robustness (Zhu et al., 2023d). When LLMs are employed as evaluators, these inherent issues can lead to suboptimal evaluation outcomes. Therefore, it is crucial to accurately and comprehensively assess the quality of LLM-as-a-judge and identify potential vulnerabilities. This section will review existing work on the evaluation of LLM-as-a-judge, focusing on three key areas: base metric (Section 4.1), bias (Section 4.2), and robustness (Section 4.3).
@@ 4.1. Basic Metric
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ is the evaluation result of LLM evaluator and human judge respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024b; Bai et al., 2023b) are also employed to access agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision, recall, and F1 scores to evaluate the performance (Zhu et al., 2023b; Wang et al., 2023d).
Both of above metrics rely on the datasets with LLM-generated response and responding human judgements. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. In (Zheng et al., 2023d), MTBench and Chatbot Arena Conversations are proposed. The former has only 80 human-crafted queries, each with several LLMs’ responses and expert-level human annotation on pairwise comparison; the latter is a large collection of crowdsourced data, with more than 30K queries from real-world users and their vote on pairs of responses from different LLMs. FairEval (Wang et al., 2023b) is based on the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated labels between ChatGPT and Vicuna responses. PandaLM (Wang et al., 2023d) constructs a test set comprising 999 pairwise samples, with queries from 252 user-oriented instructions in (Wang et al., 2022). LLMEval2 (Zhang et al., 2023d) is much larger than the previous two, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Shepherd (Wang et al., 2023c) collects 352 samples from multiple sources for its critique model as a test set to evaluate the quality of the critiques. In (Zeng et al., 2023), a meta-evaluation benchmark was created to evaluate the effectiveness of LLM evaluators in recognizing instruction-following outputs, focusing on only a part of the LLM-as-a-Judge pipeline.
More recent works propose benchmark for code (Zhao et al., 2024b), multi-modal (Chen et al., 2024b) and non-English (Son et al., 2024) tasks. Table 1 shows the benchmarks and their statistics.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023d) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023d) 2023 30k Pairwise 3 FairEval (Wang et al., 2023b) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0
@@ 4.2. Bias
Previous reviews have highlighted that large language models exhibit various types of biases across various tasks (Gallegos et al., 2024b; Dai et al., 2024; Tan et al., 2024a). These internal biases of LLMs may also affect LLM-as-a-judge, leading to unfair evaluation outcomes and subsequently impacting the development of LLMs. Therefore, it is crucial to understand the types of biases that LLM evaluators might possess and to systematically assess these biases. In this section, we systematically review various types of biases in the LLM-as-a-judge context, including their definitions, relevant metrics, and datasets that can be used for evaluation.
@@
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024b; Thakur et al., 2024; Shi et al., 2024a; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024b). To measure this bias, recent work (Shi et al., 2024a) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor response in certain positions. The study (Wang et al., 2024b) also introduced a metric Conflict Rate to measure the percent of disagreement after change the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
@@
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023d; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original response into a more verbose one (Ye et al., 2024; Zheng et al., 2023d). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, previous study (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
@@
Self-Enhancement Bias describe the phenomenon that LLM evaluators may prefer response generated by themselves (Zheng et al., 2023d; Ye et al., 2024). Considering the significant self-enhancement bias, as suggested in (Ye et al., 2024), we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
@@
Other Bias. Diversity Bias refers to bias against certain demographic groups (Ye et al., 2024), including certain genders (Chen et al., 2024a), race, and sexual orientation (Kumar et al., 2024b). As revealed in (Chen et al., 2024a), evaluators’ tendency toward visually appealing content, regardless of its actual validity, such as the text with emoji. Concreteness bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024). Furthermore, LLM evaluators may favor response with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as seyiment bias (Ye et al., 2024; Li and Sinnamon, 2023).
@@
To advance the development of LLM-as-a-Judge systems, future efforts should address two key challenges: (i) Need for Systematic Benchmark. Due to the diversity of biases, it is crucial to propose a systematic benchmark to evaluate the extent of various biases. As shown in Table 1, EVALBIASBENCH (Park et al., 2024) was proposed as a test set to measure six types of bias. Other work (Ye et al., 2024) is dedicated to proposing a unified bias testing process, including automated perturbation and a unified metric. They constructed a bias quantification framework CALM, which covers 12 types of bias. Despite these efforts, there is still no systematic benchmark and dataset that includes all types of biases. (ii) Challenges of Controlled Study. When conducting an investigation into a certain type of bias, it is challenging to isolate the specific direction of interest from other biases and quality-related characteristics. For instance, in the case of position bias, lengthening the response could potentially alter the style, fluency, and coherence, or even introduce new biases such as self-enhancement bias. Additionally, the tendency for GPT-4 to favor its own responses over those of GPT-3.5 can be interpreted as either self-enhancement bias or a proper tendency towards higher quality text. Therefore, it is essential for analytical work to carefully control for these variances.
@@ 4.3. Adversarial Robustness
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Furthermore, other work (Zheng et al., 2024) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. Other researches (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
5. Effectiveness Experiments
Figure 10. LLM-as-a-Judge Meta-evaluation Pipeline and Tools
In Section 3, We have introduced improvement strategies adopted by researchers in existing LLM-as-a-judge works to improve the evaluation capabilities of LLMs. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, as shown in Table 1, there is still a lack of meta-evaluation on whether these improvement strategies effectively optimize the LLM evaluators and which dimensions of evaluation performance are being enhanced. It is possible that some improvement strategies fail to enhance the LLM evaluators’ performance or mitigate biases in practical use, leading to computing waste. In this section, based on benchmarks mentioned in Section 4, we designed a robust and scalable meta-evaluation pipeline and tools as shown in Fig 10, and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.